# Optimal ambition in business, politics and life

### v3

@@ IV.5 Data Availability
Recall that—in the special case of a rugged landscape with Gaussian rewards—we can write down the expected cumulative reward for a given agent by summing over all the possible lengths of the exploration phase (Equation 2).
@@
$\mathbb{E}[\text{reward}]$ $=t_{\text{max}}\phi(T)\sum_{t=0}^{t=t_{\text{max}}}\Bigg{[}\Phi^{t}(T)-\Phi^{t-1}(T)\dfrac{t}{t_{\text{max}}}\Bigg{]}.$ (3)
@@
which we consider separately, where $S_{1}=\sum_{t=0}^{t=t_{\text{max}}}\Phi^{t}(T)$ is a geometric sum, and $S_{2}=\sum_{t=0}^{t=t_{\text{max}}}\Phi^{t-1}(T)\dfrac{t}{t_{\text{max}}}$ is an arithmetico-geometric sum. We can write down the partial sum expressions for both $S_{1}$:
$S_{1}$ $=\sum_{t=0}^{t=t_{\text{max}}}\Phi^{t}(T)=\dfrac{1-\Phi^{t_{\text{max}}+1}(T)}{1-\Phi(T)},$ (5)
@@
$S_{2}$ $=\sum_{t=0}^{t=t_{\text{max}}}\Phi^{t-1}(T)\dfrac{t}{t_{\text{max}}}$ (6)
@@
This result implies that given infinite time, the “explore” period is negligible compared to the “explore” stage, and the agent can be arbitrarily ambitious.
.6 Proving that the optimal threshold is greater than the landscape mean ($\mu=0$)
The reward in the “exploit” phase $\mu_{\text{exploit}}(T)$ is always positive and increases in $T$, but is close to zero for negative values of threshold $T$, so high positive thresholds yield higher rewards and negative thresholds tend to yield rewards close to the mean. The expected number of time steps it takes to satisfy threshold $T$ is $\dfrac{1}{1-\Phi(T)}$. At $T$=0, the expected length of the exploration phase is two time steps, and grows exponentially as threshold $T$ increases. Intuitively, the agents who can afford the exploration phase to be longer than two time steps should target positive thresholds $T$.
We show that the optimal threshold is above the mean for the maximally rugged landscape. We can split the summand in (2) into the reward component
@@
.6.1 The reward component increases near-linearly in $T$
As the threshold $T\rightarrow\infty$, $\mu_{\text{explore}}\rightarrow 0$, and $\mu_{\text{exploit}}\rightarrow T$. As the threshold $T\rightarrow-\infty$, $\mu_{\text{explore}}\rightarrow T$, and $\mu_{\text{exploit}}\rightarrow 0$ Small (2010). For a given $t_{x}$, as $T$ gets large, the linear combination of $\mu_{\text{explore}}$ and $\mu_{\text{exploit}}$ grows near-linearly in $T$.
.6.2 The probability component (12) is unimodal
The derivative of the probability component $P(T,t_{x})$ with respect to threshold $T$ is
@@
.6.4 Optimal threshold is greater than the mean
Since the reward component $R(T,t_{x})$ increases near-linearly in $T$, and $P(T,t_{x})$ is maximized when $\Phi(T)=\frac{t_{x}}{t_{x}+1}$, the summand $R(T,t_{x})P(T,t_{x})$ for a given $t_{x}$ achieves its maximum for a threshold $T^{*}_{t_{x}}>\Phi^{-1}\left(\frac{t_{x}}{t_{x}+1}\right)$ (See Figure 6).
As $t_{x}$ increases, $\frac{t_{x}}{t_{x}+1}$ increases also, and therefore $T^{*}_{t_{x}}$ increases, meaning that each subsequent summand achieves its maximum at a greater threshold value $T$ that the one for a previous summand.
Let $t_{x}=1$. The maximizing satisfaction threshold $T^{*}_{1}=\Phi^{-1}\left(\frac{1}{2}\right)=0.$ Since every summand $R(T,t_{x})P(T,t_{x})$ is unimodal, every summand for $t_{x}>1$ achieves its maximum for $T^{*}_{t_{x}}>0$, which implies that $\left.\frac{\partial RP(T,t_{x})}{\partial T}\right|_{T=0}>0$, $\forall t_{x}>1$. Therefore for $t_{\text{max}}>1$, the optimal threshold is greater than the mean.
Figure 6: The summands in (3) as a function of threshold $T$ and time step $t_{x}$. Each summand can be thought of as a slice for a fixed time $t_{x}$. Note that the maximum of each summand occurs at $T>0$.
Algorithm 1 Simulation algorithm
1:Input: $t_{\text{m}ax}$, $N$ agents, threshold $T$, information type, landscape mean $\mu$, landscape variance $\sigma^{2}$, smoothness parameter $\phi$, skew parameter $\alpha$.
2:Output: matrix of rewards of size $t_{\text{max}}\times N$triangleright Initialize the matrix of rewards
3:reward=matrix(nrow=$t_{\text{m}ax}$, ncol=$N$) $\triangleright$ Assign random initial rewards
4:reward[0,:]=$\epsilon_{t}\sim\mathcal{SN}(\mu,\sigma,\alpha)$, for $N$ samples. $\triangleright$ Loop through time
5:for $t$ between 1 and $t_{\text{max}}$ do $\triangleright$ Loop across agents
6:    for a in between 1 and $N$ do $\triangleright$ Compute summary statistics for all-cohort comparison
7:       if infotype==“cohort” then
8:          info=reward[t-1, !a]
9:          $\mu_{\text{info}}$=mean(info)
10:          $\sigma_{\text{info}}$=sd(info)
11:       end $\triangleright$ Compute summary statistics for upward social comparison
12:       if infotype==“upward” then
13:          info=reward[t-1, !a] $>$ reward[t-1, a]
14:          $\mu_{\text{info}}$=mean(info)
15:          $\sigma_{\text{info}}$=sd(info)
16:       end $\triangleright$ Pass summary statistics for underlying landscape
17:       if infotype==“landscape” then
18:          $\mu_{\text{info}}$=$\mu$
19:          $\sigma_{\text{info}}$=$\sigma$
20:       end $\triangleright$ Compare previous reward to threshold and move or stay
21:       prevreward=reward[t-1, a]
22:       if prevreward $<$ ($\mu+T\sigma$then
23:          newreward=$\varphi$*prevreward+
                              $(1-\varphi)\epsilon_{t}$, $\epsilon_{t}\sim\mathcal{SN}(\mu,\sigma,\alpha)$
24:       else
25:          newreward=prevreward
26:       end $\triangleright$ Update reward history
27:       reward[t,a]=newreward
28:    end for
@@
Acknowledgements
We thank members of the Burgess lab, Ellen DeGennaro, Todd Cherry, and Jonas L. Juul for comments. E.L. and M.G.B. acknowledge funding from the Cooperative Institute for Research in Environmental Sciences (CIRES) Visiting Fellows Program at the University of Colorado Boulder, funded by NOAA Cooperative Agreement NA22OAR4320151.

### v4

@@ A.1 Limiting behavior as $t_{\text{max}}\rightarrow\infty$
Recall that—in the special case of a rugged landscape with Gaussian rewards—we can write down the expected cumulative reward for a given agent by summing over all the possible lengths of the exploration phase (equation (2)).
@@
$\mathbb{E}[\text{reward}]$ $=t_{\text{max}}\phi(T)\sum_{t_{x}=0}^{t_{\text{max}}}\Bigg[\Phi^{t}(T)-\Phi^{t_{x}-1}(T)\dfrac{t_{x}}{t_{\text{max}}}\Bigg].$ (5)
@@
which we consider separately. Here, $S_{1}=\sum_{t_{x}=0}^{t_{\text{max}}}\Phi^{t}_{x}(T)$ is a geometric sum, and $S_{2}=\sum_{t_{x}=0}^{t_{\text{max}}}\Phi^{t_{x}-1}(T)\dfrac{t_{x}}{t_{\text{max}}}$ is an arithmetico-geometric sum. We can write down the partial sum expressions for both $S_{1}$:
$S_{1}$ $=\sum_{t_{x}=0}^{t_{\text{max}}}\Phi^{t}(T)=\dfrac{1-\Phi^{t_{\text{max}}+1}(T)}{1-\Phi(T)},$ (7)
@@
$S_{2}$ $=\sum_{t_{x}=0}^{t_{\text{max}}}\Phi^{t_{x}-1}(T)\dfrac{t_{x}}{t_{\text{max}}}$ (8)
@@
This result implies that given infinite time, the “explore” period is negligible compared to the “exploit” stage, and the agent can be arbitrarily ambitious.
A.2 The optimal satisfaction threshold is finite and increases as the total time increases
For any threshold, $T$, the probability of finding a strategy that satisfies the threshold increases with the length of the search, $t_{\text{max}}$. The amount of time one gets to exploit a satisfactory strategy also increases in $t_{\text{max}}$, all else equal. Both of these patterns shift incentive towards a larger optimal threshold as $t_{\text{max}}$ increases. In the limiting case of $t_{\text{max}}\rightarrow\infty$, the expected reward per time step is equal to $\mu_{\text{exploit}}$, meaning that when time is infinite, agents can be arbitrarily ambitious, but optimal ambition is finite in finite time.
A.3 The optimal satisfaction threshold is strictly greater than the mean reward
The reward in the ‘exploit’ phase ($\mu_{\text{exploit}}(T)$) is always positive and increases in $T$. The expected number of time steps it takes to satisfy threshold $T$ is $\frac{1}{1-\Phi(T)}$: at $T=0$, the expected length of the exploration phase is two time steps, and it increases exponentially as the threshold $T$ increases. Intuitively then, agents who can afford the exploration phase to be longer than two time steps should target positive thresholds $T$ (i.e., thresholds larger than the mean, $\mu=0$).
Figure 7: The summands in equation (5) as a function of threshold $T$ and time step $t_{x}$. Each summand can be thought of as a slice for a fixed time $t_{x}$. Note that the maximum of each summand occurs at $T>0$ (i.e., at a satisfaction threshold above the mean, $\mu=0$).
The reward in the “exploit” phase $\mu_{\text{exploit}}(T)$ is always positive and increases in $T$, but is close to zero for negative values of threshold $T$, so high positive thresholds yield higher rewards and negative thresholds tend to yield rewards close to the mean. The expected number of time steps it takes to satisfy threshold $T$ is $\dfrac{1}{1-\Phi(T)}$. At TT=0, the expected length of the exploration phase is two time steps, and grows exponentially as threshold $T$ increases. Intuitively, the agents who can afford the exploration phase to be longer than two time steps should target positive thresholds $T$.
We show that the optimal threshold is above the mean for the maximally rugged landscape. We can split the summand in equation (2) into the reward component
@@ A.3 The optimal satisfaction threshold is strictly greater than the mean reward
A.3.1 The reward component increases near-linearly in $T$
As the threshold $T\rightarrow\infty$, $\mu_{\text{explore}}\rightarrow 0$, and $\mu_{\text{exploit}}\rightarrow T$. As the threshold $T\rightarrow$-$\infty$, $\mu_{\text{explore}}\rightarrow T$, and $\mu_{\text{exploit}}\rightarrow 0$ [59]. For a given $t_{x}$, as $T$ gets large, the linear combination of $\mu_{\text{explore}}$ and $\mu_{\text{exploit}}$ grows near-linearly in $T$.
A.3.2 The probability component (14) is unimodal
The partial derivative of the probability component $P(T,t_{x})$ with respect to threshold $T$ is
@@ A.3.3 The case of $t_{x}=0$
A.3.4 Optimal threshold is greater than the mean
Since the reward component $R(T,t_{x})$ increases near-linearly in $T$, and $P(T,t_{x})$ is maximized when $\Phi(T)=\frac{t_{x}}{t_{x}+1}$, the summand $R(T,t_{x})P(T,t_{x})$ for a given $t_{x}$ achieves its maximum for a threshold $T^{*}_{t_{x}}>\Phi^{-1}\left(\frac{t_{x}}{t_{x}+1}\right)$ (see Fig. 7).
As $t_{x}$ increases, $\frac{t_{x}}{t_{x}+1}$ increases also, and therefore $T^{*}_{t_{x}}$ increases, meaning that each subsequent summand achieves its maximum at a greater threshold value $T$ than the one for a previous summand.
Figure 8: The optimal threshold, $T^{*}$, is closely approximated by a logarithmic function. The left panel shows the numerical estimate and the fitted logarithmic curve. The right panel shows the error $T^{*}-y$. Note that the error changes sign, meaning that the optimal threshold grows faster than the logarithmic fit.
Let $t_{x}=1$. The maximizing satisfaction threshold $T^{*}_{1}=\Phi^{-1}\left(\frac{1}{2}\right)=0.$ Since every summand $R(T,t_{x})P(T,t_{x})$ is unimodal, every summand for $t_{x}>1$ achieves its maximum for $T^{*}_{t_{x}}>0$, which implies that $\left.\frac{\partial RP(T,t_{x})}{\partial T}\right|_{T=0}>0$, $\forall t_{x}>1$. Therefore for $t_{\text{max}}>1$, the optimal threshold is greater than the mean. Fig. 8 shows that while the optimal threshold can be approximated by a logarithmic function, it grows faster than the logarithmic fit.
Appendix B Simulation Algorithm
1:Input: $t_{\max}$, NN agents, threshold $T$, information type infotype, landscape mean $\mu$, variance $\sigma^{2}$, smoothness $\varphi$, skew $\alpha$
2:Output: matrix reward of size $t_{\max}\times N$
[…]
4:reward $\leftarrow$ matrix(nrow $t_{\max}$, ncol =N=N) $\triangleright$ Assign random initial rewards
5:for $a\leftarrow 1$ to $N$ do
6:  draw $\epsilon_{0}\sim\mathcal{SN}(\mu,\sigma,\alpha)$
[…]
8:end for
[…]
10:for $t\leftarrow 1$ to $t_{\max}$ do $\triangleright$ Loop across agents
11:  for $a\leftarrow 1$ to $N$ do $\triangleright$ Summary stats per infotype
12:    if infotype == "cohort" then
13:     info $\leftarrow\{\,\texttt{reward}[t-1,j]:j\neq a\,\}$
14:     $\mu_{\text{info}}\leftarrow\texttt{mean}(\texttt{info})$,    $\sigma_{\text{info}}\leftarrow\texttt{sd}(\texttt{info})$
15:    else if infotype == "upward" then
16:     info $\leftarrow\{\,\texttt{reward}[t-1,j]:\texttt{reward}[t-1,j]>\texttt{reward}[t-1,a]\,\}$
17:     $\mu_{\text{info}}\leftarrow\texttt{mean}(\texttt{info})$,    $\sigma_{\text{info}}\leftarrow\texttt{sd}(\texttt{info})$
[…]
19:     $\mu_{\text{info}}\leftarrow\mu$,    $\sigma_{\text{info}}\leftarrow\sigma$
[…]
21:    prev $\leftarrow$ reward[t1,a][t-1,a]
22:    if prev $\mu+T\,\sigma$ then
23:     draw $\epsilon_{t}\sim\mathcal{SN}(\mu,\sigma,\alpha)$
24:     new $\leftarrow\varphi\cdot\texttt{prev}+(1-\varphi)\cdot\epsilon_{t}$
25:    else
26:     new $\leftarrow\texttt{prev}$
27:    end if
[…]
@@ Appendix B Simulation Algorithm
30:end for
Figure 9: Simulation algorithm
Appendix C Sensitivity to cohort size and cohort threshold limits in the case of social comparison
The qualitative behavior of the model in the case of social comparison is not sensitive to cohort sizes and cohort threshold limits. Fig. 10 shows the social comparison behavior for different cohort sizes. Fig. 11 shows the social comparison behavior for different threshold limit intervals. Regardless of cohort size and threshold interval, all-cohort comparison leads to higher average rewards and higher optimal thresholds than upward social comparison. Both modes of social comparison exhibit lower average reward and lower optimal threshold than the idealized reference landscape where the agents have perfect information about the mean and variance.
Figure 10: Social comparison penalizes high satisfaction thresholds, regardless of cohort size. The figure shows the mean rewards for all-cohort comparison and upward social comparison on a rough, non-skewed landscape, for cohort sizes of 10, 25, 50, 100, and 200 agents. $\epsilon_{t}\sim\mathcal{N}(\mu,\sigma^{2})$, $\varphi=0$, $t_{\text{max}}=1000$, cohorts of 100100 agents, with thresholds uniformly sampled from [3,6][-3,6], averaged over $10^{3}$ simulations.
Figure 11: Social comparison penalizes high satisfaction thresholds, regardless of threshold limits. The figure shows the mean rewards for all-cohort comparison and upward social comparison on a rough, non-skewed landscape, for a cohort size of 100 and with threshold limits samples uniformly on the intervals: [3,6][-3,6], [3,4][-3,4], [3,0][-3,0], [2,1][-2,1], [1.5,1.5][-1.5,1.5], and [1,2][-1,2]. $\epsilon_{t}\sim\mathcal{N}(\mu,\sigma^{2})$, $\varphi=0$, $t_{\text{max}}=1000$, cohorts of 100100 agents, with thresholds uniformly sampled from [3,6][-3,6], averaged over $10^{4}$ simulations.